OpenStudy (anonymous):

How to write x^2-5x+8 in the form (a-x)^2+b and hence show that x^2-5x+8>0 for all values of x ?

OpenStudy (across):

Hint:\[x^2-5x+8=x^2-5x+\frac{25}{4}+\frac{7}{4}=\left(\frac{5}{2}-x\right)^2+\frac{7}{4}\]